# ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models

### v1

@@ Abstract
The effectiveness of Large Language Models (LLMs) in solving tasks vastly depends on the quality of the instructions, which often require fine-tuning through extensive human effort. This highlights the need for automated instruction optimization; however, this optimization is particularly challenging when dealing with black-box LLMs, where model parameters and gradients remain inaccessible.
We propose ACING, a task-specific prompt optimization approach framed as a stateless continuous-action Reinforcement Learning (RL) problem, known as the continuum bandit setting. ACING leverages an actor-critic-based method to optimize prompts, learning from non-differentiable reward signals. We validate ACING by optimizing prompts for ChatGPT on 30 instruction-based tasks, as well as a summarization task. ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points compared to the best baseline considered. Furthermore, ACING not only recovers but also surpasses human-crafted expert instructions, achieving up to a 39 percentage point improvement over human benchmarks.
ACING: Actor-Critic for Instruction Learning
in Black-Box Large Language Models
[…]
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of tasks Zhao et al. (2024); Touvron et al. (2023). This success is largely attributed to their strong instruction-following capabilities, which enable adaptation to diverse downstream applications Chen et al. (2023); Liu et al. (2023). These instructions, commonly referred to as prompts, play a crucial role in the performance of LLMs Wei et al. (2022); Zhu et al. (2024); Liu et al. (2023). Given the importance of prompts, researchers are increasingly interested in automating their optimization to reduce the need for manual adjustments, which is often a labor-intensive and costly process Reynolds and McDonell (2021); Mishra et al. (2021). As a result, developing efficient methods for automatic prompt optimization has become highly important to maximize LLMs performance.
Several methods for automated prompt optimization have been proposed in the literature. Some focus on optimizing continuous prompts, known as soft prompts Zhong et al. (2021); Li and Liang (2021), which are typically fed into a LLM after its embedding layer since they do not represent meaningful language that the LLM can interpret directly. Other methods aim at optimizing discrete prompts, known as hard prompts, by either generating diverse prompts Zhou et al. (2023) or refining existing ones Pryzant et al. (2023). For both soft and hard prompts, some approaches rely on gradient-based methods Shin et al. (2020); Lester et al. (2021); Li and Liang (2021). However, these methods require access to the gradients of the LLMs and are therefore limited to white-box LLMs. Meanwhile, the most powerful LLMs today are typically black-box models (e.g., ChatGPT OpenAI (2023a) and GPT-4 OpenAI (2023b)).
To overcome these challenges, recent studies have introduced gradient-free techniques for optimizing prompts in black-box LLMs Zhou et al. (2023); Prasad et al. (2023); Pryzant et al. (2023). While promising, these approaches rely on heuristic local search methods and fail to fully utilize the observation history, including previously tested instructions and their outcomes, when selecting new instructions to test. Consequently, they are less effective at balancing the inherent exploration-exploitation trade-off, leading to inefficient querying. This inefficiency becomes problematic and impractical when API calls to black-box LLMs involve significant time and financial costs.
A recent line of work has focused on optimizing instructions for black-box LLMs by searching through soft prompts and transforming them into discrete prompts using white-box LLMs. InstructZero Chen et al. (2024) optimizes discrete prompts for ChatGPT by applying Bayesian optimization over soft prompts. Building on the same line, INSTINCT Lin et al. (2024) replaces Bayesian optimization with NeuralUCB Zhou et al. (2020), a critic algorithm, to select actions, from a finite pool of soft prompts, initially generated with Sobol Eriksson et al. (2019).
Inspired by the success of actor-critic methods for control problems Haarnoja et al. (2018, 2019) and generative adversarial networks Goodfellow et al. (2014), our approach combines the hybrid white-box/black-box framework with an actor-critic Reinforcement Learning (RL) algorithm, operating in a stateless, infinite continuous-action space, and constrained by a strict budget of API calls. The actor aims to maximize the critic’s value while remaining as stochastic as possible to enhance exploration and simultaneously exploit promising actions. This approach increases the likelihood of finding optimal prompts and enables more efficient use of the limited number of API calls. ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points. Furthermore, our optimized prompts not only recover but also surpass human-crafted expert instructions, achieving up to a 39 percentage point improvement against human benchmarks.
An implementation of ACING is available at https://github.com/salmakh1/ACING.
@@ 2 Problem Formulation
2.1 Prompt Learning Objective
We aim to improve the performance of a black-box LLM, denoted by $f$, which can only be accessed through its API, while its internal parameters remain unknown. Given a task represented by an (unknown) distribution $(x,y)\sim\mathcal{D}$—where $x$ denotes possible inputs and $y$ the corresponding correct outputs—our goal is to find the optimal prompt $\tau^{\star}$ that maximizes the likelihood of $f$ producing correct outputs for a given task. This is evaluated using a scoring function $q(\cdot,\cdot)\in[0,1]$. The black-box model $f$ processes an input formed by concatenating ($\oplus$) the prompt $\tau$ with the sentence $x$, producing a predicted output $\hat{y}=f(\tau\oplus x)$. More formally, the objective is to maximize the expected score of the LLM in solving the task $\mathcal{D}$, defined as:
@@ 2.1 Prompt Learning Objective
To achieve this, we utilize a validation dataset $\mathcal{V}=\{(x_{j},y_{j})\}_{j=1}^{m}$, where each pair consists of an input sentence $x_{j}$ and its corresponding ground truth output $y_{j}$. Our objective is to find the prompt that maximizes the scoring function $q(\cdot,\cdot)$ across the validation dataset, where $q(\hat{y}_{j},y_{j})$ measures the quality of the predicted output $\hat{y}_{j}$ against the true output $y_{j}$. Thus, our objective becomes finding the prompt $\tau^{\star}$ that maximizes the average score over the validation set $\mathcal{V}$. The derived prompt $\tau^{\star}$ is then evaluated on a separate test set $\mathcal{T}=\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$ to assess its generalization performance.
2.2 Learning over a Continuous Space
Directly optimizing the prompt $\tau$ for the black-box LLM model $f$ presents substantial challenges due to the discrete combinatorial nature of token selection in $\tau$. To mitigate this challenge, prior approaches like InstructZero Chen et al. (2024) and INSTINCT Lin et al. (2024) employ a white-box LLM, represented by $h$, to convert the discrete optimization problem into a continuous one by introducing a soft prompt vector $\mathbf{z}\in\mathbb{R}^{d}$, which is a continuous $d$-dimensional vector representing the token embedding of a set of virtual tokens.
The white-box model $h$ has access to a dataset of exemplars, =$\mathcal{E}=\{(u_{j},v_{j})\}_{j=1}^{k}$, where each pair $(u_{j},v_{j})$ defines input-output text sequences that exemplify a downstream task. The white-box LLM $h$ serves as a proxy, mapping the soft prompt $\mathbf{z}$ into a discrete prompt $\tau$ suitable for input to the black-box LLM. Specifically, given the exemplars $\mathcal{E}$ and the vector $\mathbf{z}$, their concatenation is input to the white-box LLM, generating the discrete prompt $\tau(\mathbf{z})=h(\mathbf{z},\mathcal{E})$. This generated prompt $\tau(\mathbf{z})$ is prepended to a test input $x_{j}$ from the validation set $\mathcal{V}$, and the combined input is provided to the black-box LLM $f$ to generate an output $\hat{y}_{j}=f(\tau(\mathbf{z})\oplus x_{j})$. The output $\hat{y}_{j}$ is then evaluated against the true output $y_{j}$ using the scoring function $q(\hat{y}_{j},y_{j})$. By using a fixed set of exemplars $\mathcal{E}$, the original discrete problem (Equation (1)) of finding the optimal prompt $\tau$ is effectively transformed into a continuous optimization problem over the soft prompt vector $\mathbf{z}$, as follows:
@@ 2.2 Learning over a Continuous Space
2.3 Reducing the Continuous Space
The soft prompt $\mathbf{z}$ is typically high-dimensional (e.g., $d=5120\times N_{z}$, with $N_{z}$ the number of virtual tokens, when $h$ is Vicuna-13B Chiang et al. (2023)), which makes it a challenging problem. Therefore, we employ random projection techniques to reduce the input dimension as done in prior works Chen et al. (2024); Lin et al. (2024). Specifically, given a matrix $P\in\mathbb{R}^{d\times d^{\prime}}$ where $d^{\prime}\ll d$, with randomly sampled elements, and a continuous vector $\mathbf{a}\in\mathbb{R}^{d^{\prime}}$, the vector $\mathbf{z}=P\mathbf{a}$ is used as the soft prompt. By substituting $\mathbf{z}$ with $P\mathbf{a}$, the input variable to be optimized in Eq. 2 becomes $\mathbf{a}$, thereby reducing the input dimension of the optimization problem to $d^{\prime}$:
@@ 2.3 Reducing the Continuous Space
For fairness with previous works, the reduced input dimension $d^{\prime}$, referred to as the intrinsic dimension, is chosen as $d^{\prime}=10$ Chen et al. (2024); Lin et al. (2024). An ablation study on the intrinsic dimension is discussed in § 5.4.3. Furthermore, to maintain consistency with these prior studies, we similarly use a projection matrix $P$ with elements uniformly and independently sampled from $\text{Uniform}(-1,1)$, and bound the vector such that $\mathbf{a}\in[0,1]^{d^{\prime}}$.
3 Actor-Critic Approach for Instruction Learning in Black-Box LLMs
3.1 Instruction Learning as an RL Problem
In this work, we frame the prompt-learning problem as an agent navigating an infinite action space within a $d^{{}^{\prime}}$-dimensional unit hypercube, searching within an infinite number of actions $\mathbf{a}\in[0,1]^{d^{{}^{\prime}}}$, in contrast to previous works that consider a finite number of actions initially generated with Sobol Eriksson et al. (2019); Lin et al. (2024). As part of the environment, these continuous actions are projected into a higher-dimensional space (§ 2.3) and then transformed into discrete prompts, through a white-box LLM (§ 4). These discrete prompts are evaluated by a black-box LLM over a validation dataset, after which a reward is computed and returned to the agent. We consider a finite horizon $T$, representing the maximum number of actions played (i.e., the budget of evaluated instructions), after which the agent ceases exploration. The agent’s goal is to discover the optimal continuous action, $\mathbf{a}^{\star}$, which produces the optimal discrete prompt and the highest reward.
Unlike in general RL settings, where the agent encounters multiple states with varying reward distributions, the agent we consider operates in a stateless, stochastic environment, commonly referred to as a stochastic bandit setting Slivkins (2019); Lattimore and Szepesvári (2020). In this context, at each iteration $t\leq T$ (where $t$ is incremented with every black-box LLM call), the agent’s action $\mathbf{a}_{t}$ generate stochastic rewards, $r_{t}\in[0,1]$, without affecting the reward distributions associated with those actions. The prompt-learning objective is modeled as a stochastic best-arm identification problem in a stateless, continuous action space RL, specifically in an infinite continuum bandit setting.
Algorithm 1 Stateless Soft Actor-Critic
0:  Evaluation budget $T$
1:  Initialize $\boldsymbol{\theta}$, $\mathbf{w}$ (e.g., randomly)
2:  for $t\leftarrow 1$ to $T$ do
3:     Choose action $\mathbf{a}_{t}\sim\pi(\cdot\mid\boldsymbol{\theta})$
4:     Take action $\mathbf{a}_{t}$ and observe reward $r_{t}$
5:     $\mathbf{w}\leftarrow\mathbf{w}-\lambda\hat{\nabla}\mathbf{J}_{Q}(\mathbf{w})\quad$ (Equation (5))
6:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\beta\hat{\nabla}\mathbf{J}_{\pi}(\boldsymbol{\theta})\quad$ (Equation (6))
7:     $\alpha\leftarrow\alpha-\gamma\hat{\nabla}\mathbf{J}_{\alpha}(\alpha)\quad$ (Equation (7))
8:  end for
3.2 Actor-Critic Continuum Bandits
RL techniques are generally used to learn optimal policies and actions leading to best rewards. These techniques can be classified into value-based (critic-centric), policy-gradient (actor-centric), or hybrid approaches. Actor-critic algorithms merge the benefits of both by employing two key components: the actor, responsible for policy optimization, and the critic, which evaluates the policy (Sutton and Barto, 2018). This hybrid approach has demonstrated success in various domains, including continuous control problems Mnih et al. (2016); Haarnoja et al. (2018, 2019) and with generative adversarial networks (GANs) (Goodfellow et al., 2014), where a generator and a critic network collaboratively optimize performance.
Figure 1: Pipeline of ACING. In each iteration, a soft prompt along with several examples of the target task are provided to the white-box LLM to generate an instruction. This instruction is then used to query the black-box LLM, which produces answers to the target task queries. The resulting score is returned to the agent as a reward, which is used to update its networks and adjust its policy. Both LLMs remain frozen throughout the process.
Inspired by the success of actor-critic approaches, we introduce a (stateless) actor-critic algorithm, as provided in Algorithm 1, tailored to our infinite continuum bandit setting, enabling autonomous learning of effective prompts with a constrained evaluation budget and outperforming previous state-of-the-art black-box prompt learning approaches. Both the actor and the critic are parameterized by neural networks. In typical RL problems, the agent interacts with an environment with varying states, so its policy depends on the current state. That is, the policy network generates actions based on the current state $s$, $\pi(.|s;\theta)$, where $\theta$ denotes the parameters. However, in the problem we consider, the setting is stateless. Therefore, we define a policy network $\pi(.;\theta)$ with a constant input that produces a $d^{{}^{\prime}}$-dimensional vector as output $\mathbf{a}$, representing the continuous action. For the critic, we use a neural network with a $d^{{}^{\prime}}$-dimensional input (the action) and a single output that estimates the action’s quality, thereby assessing the quality of the policy that proposed it. The critic network’s parameters are trained to minimize the following:

### v2

# ACING: Actor-Critic for Instruction Learning in Black-Box LLMs

@@ Abstract
The effectiveness of Large Language Models (LLMs) in solving tasks depends significantly on the quality of their instructions, which often require substantial human effort to craft. This underscores the need for automated instruction optimization. However, optimizing instructions is particularly challenging when working with black-box LLMs, where model parameters and gradients are inaccessible. We introduce ACING, an actor-critic reinforcement learning framework that formulates instruction optimization as a stateless, continuous-action problem, enabling exploration of infinite instruction spaces using only black-box feedback. ACING automatically discovers prompts that outperform human-written prompts in 76% of instruction-induction tasks, with gains of up to 33 points and a 10-point median improvement over the best automatic baseline in 33 tasks spanning instruction-induction, summarization, and chain-of-thought reasoning. Extensive ablations highlight its robustness and efficiency. An implementation of ACING is available at https://github.com/salmakh1/ACING.
ACING: Actor-Critic for Instruction Learning in Black-Box LLMs
Salma Kharrat          Fares Fourati KAUST {salma.kharrat, fares.fourati}@kaust.edu.sa          Marco Canini
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across tasks like classification, summarization, and reasoning Cobbe et al. (2021); Touvron et al. (2023); Zhao et al. (2024). A key driver of this success is their ability to follow natural language instructions, commonly known as prompts Chen et al. (2023a); Wei et al. (2022); Liu et al. (2023). Yet crafting effective prompts remains labor-intensive and brittle, particularly in black-box settings where model internals are inaccessible.
A core challenge in prompt engineering lies in the extreme sensitivity of LLMs to subtle linguistic variations. Even minor changes in phrasing can lead to substantial performance shifts. As shown in Table˜1, a minor wording change yields a 12-point gain (using GPT-4o Hurst et al. (2024)) despite preserving semantics. Such prompt sensitivity is pervasive, underscoring the need for robust, automated prompt optimization methods.
Recent work has explored automating prompt design to reduce human effort Reynolds and McDonell (2021); Mishra et al. (2021). Soft prompting techniques Li and Liang (2021); Lester et al. (2021) and heuristic-based discrete search methods Zhou et al. (2023); Pryzant et al. (2023) have shown promise, yet each faces key limitations in black-box settings. Soft prompts require access to the model internals, limiting them to white-box scenarios. Discrete search methods, in turn, often struggle to explore vast and nuanced instruction spaces efficiently. More recent hybrid methods combine white-box prompt generation with black-box evaluation Chen et al. (2023b); Lin et al. (2024b), but typically rely on finite candidate pools or rigid reward assumptions, constraining their applicability.
To overcome these limitations, we introduce ACING, an actor-critic reinforcement learning (RL) framework for automated instruction optimization in black-box LLM settings. ACING formulates prompt optimization as a stateless, continuous-action RL problem within a continuum bandit environment. By learning a latent instruction space through an off-policy actor-critic algorithm, ACING efficiently explores infinite instruction candidates, using only black-box feedback for evaluation. A frozen white-box model is used as a decoder to convert latent vectors into discrete prompts, enabling both scalability and linguistic richness. The generated instructions are not only high-performing but also naturally interpretable, clear, and semantically aligned with the tasks, making them suitable for practical deployment. To effectively guide exploration under tight query budgets, ACING leverages entropy-regularized policy optimization, encouraging diversity in sampled prompts and improving the likelihood of discovering high-performing instructions.
Source Instruction Score
Human Write a word that means the opposite of the input word. 0.70
ACING Take a word and change it to its opposite. 0.82
Table 1: Our prompt vs human on GPT-4o for antonym task.
Notably, to our knowledge, ACING is the first approach to apply off-policy, continuous-action actor-critic reinforcement learning to instruction learning in black-box LLMs, using lightweight actor and critic neural networks (rather than LLMs), making the approach both efficient and widely applicable.
We evaluate ACING against state-of-the-art techniques, including Bayesian optimization, contextual bandits, and evolutionary strategies. Empirical results across 33 diverse tasks—including instruction induction, reasoning (including zero-shot Chain-of-Thought (CoT)), semantics, syntax, phonetics, translation, summarization, and code understanding—show that ACING outperforms both human-written prompts and strong automated baselines. Notably, it surpasses human instructions in 76% of instruction induction tasks, achieving gains of up to 33 points and a median improvement of 10 points over the best automatic baselines.
In summary, our contributions are threefold:
(1) An RL formulation of instruction learning: We propose a continuous-action actor-critic RL approach for instruction optimization in black-box LLMs, enabling scalable exploration of infinite instruction spaces using entropy-regularized policies.
(2) Comprehensive validation. Across 33 diverse tasks, including instruction induction, zero-shot CoT reasoning, and summarization, ACING consistently outperforms both human-written and strong automated instructions, achieving statistically significant gains.
(3) In-depth analysis and insights: Through extensive ablation studies, we analyze the impact of latent dimensionality, exemplar configuration, decoder architecture, and optimization budget, highlighting the robustness of the ACING approach. Additionally, human evaluation and automated readability analysis confirm the clarity and semantic faithfulness of the generated instructions.
Figure 1: Pipeline of ACING. At each iteration, a soft prompt and task exemplars are fed to the white-box model to generate an instruction. This instruction queries the black-box LLM, whose outputs are scored. The resulting score is returned to the agent as a reward, which is used to update its networks and adjust its policy. Both LLMs remain frozen throughout.
@@ 2 Problem Formulation
2.1 Problem: Prompt Optimization for Black-Box LLMs
We aim to improve the performance of a black-box LLM, denoted by $f$, which can only be accessed through its API, while its internal parameters remain unknown. Given a task represented by an (unknown) distribution $(x,y)\sim\mathcal{D}$—where $x$ denotes possible inputs and $y$ the corresponding correct outputs—our goal is to find the optimal prompt $\tau^{\star}$ that maximizes the likelihood of $f$ producing correct outputs for a given task. This is evaluated using a scoring function $q(\cdot,\cdot)\in[0,1]$.
The black-box model $f$ processes an input formed by concatenating ($\oplus$) the prompt $\tau$ with the sentence $x$, producing a predicted output $\hat{y}=f(\tau\oplus x)$. More formally, the objective is to maximize the expected score of the LLM in solving the task represented by the distribution $\mathcal{D}$, defined as:
@@ 2.1 Problem: Prompt Optimization for Black-Box LLMs
We utilize a validation dataset $\mathcal{V}=\{(x_{j},y_{j})\}_{j=1}^{m}$, where each pair consists of an input sentence $x_{j}$ and its corresponding ground truth output $y_{j}$. Our objective is to find the prompt that maximizes the scoring function $q(\cdot,\cdot)$ across the validation dataset, where $q(\hat{y}_{j},y_{j})$ measures the quality of the predicted output $\hat{y}_{j}$ against the true output $y_{j}$. Thus, our objective becomes finding the prompt $\tau^{\star}$ that maximizes the average score over the validation set $\mathcal{V}$. The derived prompt $\tau^{\star}$ is then evaluated on a separate test set $\mathcal{T}=\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$ to assess its generalization performance.
2.2 Reformulating Discrete Prompt Search as a Continuous Optimization Problem
Directly optimizing the prompt $\tau$ for the black-box LLM model $f$ presents substantial challenges due to the discrete combinatorial nature of token selection in $\tau$. To mitigate this challenge, similar to prior approaches Chen et al. (2024); Lin et al. (2024b); Hu et al. (2024) we employ a publicly available, open-source white-box model, represented by $h$, and introduce a soft prompt vector $\mathbf{z}\in\mathbb{R}^{d}$, which is a continuous $d$-dimensional vector representing the token embedding of a set of virtual tokens. The white-box model $h$, which remains entirely frozen, with no training or gradient updates, serves as a proxy mapping $\mathbf{z}$ into a discrete prompt $\tau$ for the black-box LLM.
Given a dataset of exemplars, $\mathcal{E}=\{(u_{j},v_{j})\}_{j=1}^{k}$, where each pair $(u_{j},v_{j})$ defines input-output text sequences that exemplify a downstream task and the vector $\mathbf{z}$, their concatenation is input to the white-box model, generating the discrete prompt $\tau(\mathbf{z})=h(\mathbf{z},\mathcal{E})$. This generated prompt $\tau(\mathbf{z})$ is prepended to a test input $x_{j}$ from the validation set $\mathcal{V}$, and the combined input is provided to the black-box LLM $f$ to generate an output $\hat{y}_{j}=f(\tau(\mathbf{z})\oplus x_{j})$. The output $\hat{y}_{j}$ is then evaluated using the scoring function $q(\hat{y}_{j},y_{j})$. By using a fixed set of exemplars $\mathcal{E}$, the original discrete problem (Eq. (1)) of finding the optimal prompt $\tau$ is effectively transformed into a continuous optimization problem over the soft prompt vector $\mathbf{z}$, as follows:
@@ 2.2 Reformulating Discrete Prompt Search as a Continuous Optimization Problem
The soft prompt $\mathbf{z}$ is typically high-dimensional. Therefore, we employ random projection techniques to reduce the input dimension as done in prior works Chen et al. (2024); Lin et al. (2024b). Specifically, we sample a matrix $P\in\mathbb{R}^{d\times d^{\prime}}$ with entries from $\text{Uniform}(-1,1)$, and optimize a lower-dimensional vector $\mathbf{a}\in[0,1]^{d^{\prime}}$. The soft prompt is then given by $\mathbf{z}=P\mathbf{a}$, transforming the original problem into optimization over a compact and continuous space as follows:
@@
This optimization problem is central to our framework. The following section elaborates on our approach to solving it. For clarity, we summarize all symbols used in our actor-critic framework in Table 6.
Figure 2: Illustration of the prompt generation and testing inside the environment using the larger_animal dataset as an example.
3 Framework for Instruction Learning
We formulate the problem of prompt learning for black-box LLMs as a RL problem, where the agent explores an infinite instruction space by sampling continuous actions $\mathbf{a}\in[0,1]^{d^{{}^{\prime}}}$. Each action corresponds to a latent prompt representation, which is mapped—via a fixed projection matrix and a decoder—into a discrete instruction. This instruction is then evaluated by the black-box LLM on a validation set to produce a reward indicating task performance. The stateless, stochastic setup places the problem in the continuum bandit regime, unlike traditional discrete multi-armed bandits Slivkins (2019); Lattimore and Szepesvári (2020).
3.1 Actor-Critic for Instruction Optimization
To efficiently learn in a high-dimensional and limited-feedback setting, we propose a stateless, off-policy actor-critic framework inspired by recent advances in continuous control Konda and Tsitsiklis (1999); Mnih et al. (2016); Haarnoja et al. (2018, 2019). Our framework comprises lightweight networks: (1) a policy network (actor) $\pi(\cdot;\theta)$, which outputs latent action vectors corresponding to soft prompts, and (2) two critics to compute a value function $Q(\mathbf{a})=\min\{Q_{\mathbf{w_{1}}}(\mathbf{a}),Q_{\mathbf{w_{2}}}(\mathbf{a})\}$, which estimates the expected reward of a given action. Formally, the policy $\pi(\cdot;\theta)$ is modeled as a Gaussian distribution with a diagonal covariance matrix. The actor network outputs the mean and log standard deviation for each action dimension, and actions are sampled using the reparameterization trick. The critics, $Q_{\mathbf{w_{1}}}(.)$ and $Q_{\mathbf{w_{2}}}(.)$, are trained to minimize the mean squared error between predicted and observed rewards: